OpenStudy (anonymous):

Let f(x) 12/4x+2= Find f(-1). a. -6 b. -2 c. 2 d. 6

OpenStudy (anonymous):

in gunna say that its c btw will u plez message me

OpenStudy (anonymous):

\[f(x)=\frac{12}{4x+2}\] \[f(-1)=\frac{12}{4\times (-1)+2}\]

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

im realy not the best at math ether but i think that it is c

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@satellite73